OpenStudy (anonymous):

Use Descartes' Rule of Signs to determine the possible numbers of positive and negative real zeros of f. Then, list the possible rational zeros of f and graph f so that some of the possible zeros in parts A and B can be disregarded and determine all of the real zeros of f. The function given is f(x)=-3x^3+20x^2-36x+16

OpenStudy (anonymous):

do you really have to do all that or can you just cut to the chase and list the three zero?

OpenStudy (anonymous):

I have to do all of those steps :( lol

OpenStudy (anonymous):

do you think you could still help me out? this is my last problem and I would realllly appreciate it

OpenStudy (anonymous):

is there maybe a typo in your question? check carefully, because there are two complex zeros and one irrational zero of this one

OpenStudy (anonymous):

oh maybe i did not see the minus sign in front!

OpenStudy (anonymous):

It's from a textbook so i dont know :o I'll type exactly what it says! "In Exercises 61-66, (a) use Descartes's Rule of Signs to determine the possible numbers of positive and negative real zeros of f, (b) list the possible rational zeros of f, (c) use a graphing utility to graph f so that some of the possible zeros in parts (a) and (b) can be disregarded, and (d) determine all the real zeros of f."

OpenStudy (anonymous):

and then the one that I was assigned to work on was 62. f(x)=-3x^3+20x^2-36x+16

OpenStudy (anonymous):

ok fine lets do d first \(-(3 x-2) (x-2) (x-4)\) is what you have so the zeros are \[2,4,\frac{2}{3}\]

OpenStudy (anonymous):

now to A use Descartes's Rule of Signs to determine the possible numbers of positive and negative real zeros of f, here \[f(x)=-3x^3+20x^2-36x+16\] and there are three changes in sign of the coefficients, from \(-3\) to \(+20\) then from \(+20\) to \(-36\) and then again from \(-36\) to \(+16\)

OpenStudy (anonymous):

that means there could be 3 positive zeros, and you count down by twos, so 3 positive zeros, or 1 positive zero

OpenStudy (anonymous):

for the negative zeros we find the changes in sign of \(f(-x)\) and \[f(-x)=3x^3+20x^2+36x+16\] this has no changes in sign so there are no negative zeros for this function

OpenStudy (anonymous):

did i lose you yet?

OpenStudy (anonymous):

not yet :p this is starting to make some sense!

OpenStudy (anonymous):

i hope it is clear how i computed \(f(-x)\) easily i just changed the sign of degree 3 and 1 term, left the others alone so to summarize there are either 3 or 1 positive zero, no negative zeros that means either 3 positive zeros or 1 positive zero and 2 complex zeros

OpenStudy (anonymous):

because since it has degree 3, there are three zeros all together, not necessarily all real now on to part B

OpenStudy (anonymous):

, (b) list the possible rational zeros of f, similar to the last one all numbers that divide 16 or fractions where the numerator divides 16 and the denominator divides the leading coefficient of 3

OpenStudy (anonymous):

like last time you get \[\pm1,\pm2,\pm4,\pm8,\pm16\] but not also \[\pm\frac{1}{3},\pm\frac{2}{3},\pm\frac{4}{3},\pm\frac{8}{3},\pm\frac{16}{3}\] quite a long list

OpenStudy (anonymous):

+/-1, +/-2, +/-4, +/-8, +/-16 and then all of that over 3?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now since we know there are no negative zeros, we can cut the list in half and get rid of the negative ones

OpenStudy (anonymous):

(c) use a graphing utility to graph f so that some of the possible zeros in parts ok lets do it, we should have done this at the start http://www.wolframalpha.com/input/?i=-3x^3%2B20x^2-36x%2B16

OpenStudy (anonymous):

so just 1,2,4,8,16,1/3,2/3,4/3,8/3, and 16/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now check the link i sent you above you get the graph in it

OpenStudy (anonymous):

oh wow and thats it?

OpenStudy (anonymous):

you see there are 3 positive zeros, and they are between 0 and 4, and in fact they are \(\frac{2}{3},2,4\)

OpenStudy (anonymous):

which means this one factors as \[-(3 x-2) (x-2) (x-4)\]

OpenStudy (anonymous):

and yes, that is it!

OpenStudy (anonymous):

wow thank you SO much!!! I really appreciate it!!! :)

OpenStudy (anonymous):

if i were to do this same exact problem with all parts but a different function of f(x)=4x^4-17x^2+4, how would I begin this one? I know that there are 2 sign changes but I dont know how to get the rest just yet

OpenStudy (anonymous):

so far I have that there are 2 sign changes which means 2 positive zeros or 0 positive zeros then, I did f(-x) to find negative zeros and got 4x^4+17x^2+4

OpenStudy (anonymous):

there are no sign changes which means no negative zeros right?

OpenStudy (anonymous):

then i found factors of 4 which were +/-1,+/-2,+/-4 but since there are no negative zeros it is only 1,2,4

OpenStudy (anonymous):

am i right so far? @satellite73

OpenStudy (anonymous):

hold on a moment

OpenStudy (anonymous):

\[f(x)=4x^4-17x^2+4\] two changes in sign so either 2 positive zeros or no positive zeros

OpenStudy (anonymous):

but \(f(-x)\) is not right all the exponents are even, so \(f(-x)=4x^7-17x^2+4\) same as \(f(x)\) the function is "even"

OpenStudy (anonymous):

this means 2 or no negative zeros as well

OpenStudy (anonymous):

hm.. i factored the function and got (x-2)(x+2)(2x-1)(2x+1) which gave me x=2 x=-2 x=1/2 x=-1/2

OpenStudy (anonymous):

oh man! so did that just mess up all of the math that i did?

OpenStudy (anonymous):

no no you are right it is all right

OpenStudy (anonymous):

so would my final answer be x=2 x=-2 x=1/2 x=-1/2 AND 2 positive, 2 negative zeros?

OpenStudy (anonymous):

you only wrote one thing wrong \[f(-x) =4x^4+17x^2+4\] that is NOT correct

OpenStudy (anonymous):

it should be \(f(-x)=4x^4-17x^2+4\) the rest is right

OpenStudy (anonymous):

and your zeros are right also

OpenStudy (anonymous):

\[-2,2,-\frac{1}{2},\frac{1}{2}\] all that is correct two positive, two negative

OpenStudy (anonymous):

oh i see thats what i meant to type on here but i have it right on my paper :p sorry about that! thank you so much!!! :)

OpenStudy (anonymous):

yw, hope this helped and it makes some sense

OpenStudy (anonymous):

yes it really did thats how i was able to do this problem! many thanks to you, you are awesome!!! :)

OpenStudy (anonymous):

yw, my pleasure